Mathematics 242 – Principles of Analysis Definitions Quiz 1 – January 31, 2014

Your Name:

A) (5) What is the definition of the intersection of two sets: A ∩ B?

B) (5) Let f : {A, B, C, D} → {1, 2, 3, 4, 5} be a function. Is it possible for f to be onto? Why or why not?